Abe is twice as old as Connie, and Bert is five years older than Connie. The sum of their ages is 30. How old are they?
Answer by Boreal(15235)

Connie is x years old
Abe is 2x
Bert is x+5
They sum to 30
x+2x+x+5=30.
Collect terms.
4x+5=30
subtract 5 from each side
4x=25
Divide both sides by 4
x=6.25
2x=12.50
x+5=11.25
Connie is 6 1/4 years old or 6.25
Abe is 12 1/2 years old or 12.5
Bert is 11 1/4 years old or 11.25
They sum to 30.